OpenStudy (anonymous):

jawad's weekly salary is $70 less than pasha,s which is $50 more than maria's.if maria's earn $280 per week,how much does jawad earn per week?

OpenStudy (anonymous):

$70

OpenStudy (hoblos):

maria's salary is $280 => pasha,s salary is $280+$50=$330 =>awad's weekly salary is 330-70=$260

OpenStudy (anonymous):

J,P and M are the people's salary. J=P-70 P=50+M M=280 P=50+280=330 J=330-70=$260